Find the expression of a polynomial function of degree 3 with the following characteristics: zeros:3,i,-i, goes through point (1,-4) please help
Answer by tommyt3rd(5050)

p(x)=a(x-3)(x-i)(x+i)
p(x)=a(x-3)(x^2+1)
p(1)=-4 means
-4=a(1-3)(1^2+1)
-4=a(-2)(2)
a=1
p(x)=1(x^3-3x^2+x-3)
p(x)=x^3-3x^2+x-3

:)